Solve the problems below and show the complete solution.

1.) Calculate the pH of a carbonate buffer system that contains 0.0065 M HCO3- and 0.0084 M CO32-.
( Ka of HCO3- is 4.7x10-11)
1
Expert's answer
2020-12-06T13:53:59-0500

HCO3- = H+ + CO32-

Ka=[H+]*[CO32-]/[HCO3-] = [H+]*c(CO32-)/c(HCO3-) = 4.7*10-11

[H+] = Ka*c(HCO3-)/c(CO32-) = 4.7*10-11*0.0065/0.0084 = 3.6369*10-11

pH = -log10([H+]) = -log10(3.6369*10-11) = 10.44